Les Hinckle made a deposit in the bank on January 1, 2005. The banks pays interest at the rate of 8% compounded annually. On January 1, 2012, the deposit has accumulated to $30,000. How much money did Les originally deposit on January 1, 2005?
Answer by lwsshak3(11628)

Les Hinckle made a deposit in the bank on January 1, 2005. The banks pays interest at the rate of 8% compounded annually. On January 1, 2012, the deposit has accumulated to $30,000. How much money did Les originally deposit on January 1, 2005?
***
Compound interest formula: A=P(1+i)^n, P=initial investment, i=interest per compounding period, n=number of periods, A=amount after n-periods.
For given problem:
i=8%=.08
n=7
A=$30,000
..
A=P(1+i)^n
30000=P(1+.08)^7=P(1.08)^7
P=30000/(1.08)^7=17505
How much money did Les originally deposit on January 1, 2005? $17,505